# Positivity is undecidable in tensor products of free algebras

### 09 Dec 2023 v1

@@ 4.2. Construction of $\mathcal{A}_{\mathcal{S}}(m)$
and similarly $\widetilde{\pi}(\widetilde{Z})\widetilde{\pi}(Q)=\widetilde{\pi}(Q)\widetilde{\pi}(\widetilde{Z})=\widetilde{\pi}(Z_{m,0})\widetilde{\pi}(Q)$. For (R4), recall from the presentation of $G_{\mathcal{S}}$ that $\pi(X_{m,0})=\pi(X_{m,n+1})$ and $\pi(S)\pi(X_{m,i})\pi(S^{*})=\pi(X_{m,i+1})$ for all $i\in\mathbb{Z}$. Since $CE_{i,i}C^{-1}=E_{i+1,i+1}$ for all $i\in\mathbb{Z}_{n+1}$,
@@
and $P_{n+1}:=P_{0}$. Recall from the presentation of $G_{\mathcal{S}}$ that $J$ is a unitary of order 2 and is central. Hence in $\mathbb{C}H_{\mathcal{S}}$, $\tfrac{1-J}{2}$ and $\tfrac{1+J}{2}$ are projections, and $(\tfrac{1+J}{2})P_{i}=0$ for $0\leq i\leq n$. Also for all $10\leq i\leq n-$ 1, $X_{mi}Z_{mi}X_{mi}Z_{mi}=J$, $X_{mi}$ commutes with $P_{i+1}$, and $P_{i}=(\tfrac{1-Z_{mi}}{2})P_{i+1}$. Hence $21+JX_{mi}Z_{mi}X_{mi}Z_{mi}=$ 2, and
@@
It follows that $\pi\big{(}(P_{i}+X_{mi}P_{i}X_{mi}-P_{i+1})(1+JX_{mi}Z_{mi}X_{mi}Z_{mi})\big{)}=0$ for all $0\leq i\leq n$, and hence $A_{1}=A_{2}=0$. We conclude that $\widetilde{\pi}(r)=0$ for all $r$ in (R0)-(R6). Hence $\widetilde{\pi}$ induces a representation of $\mathcal{A}_{\mathcal{S}}(m)$ on $\widetilde{\mathcal{H}}\oplus\widetilde{\mathcal{H}}$.
The image of $\widetilde{\pi}$ is contained in $M_{2}\big{(}\pi(\mathbb{C}H_{\mathcal{S}})\otimes\mathcal{B}(\mathcal{H}_{1})\big{)}$ This algebra is isomorphic to $\pi(\mathbb{C}H_{\mathcal{S}})\otimes M_{2(n+1)}(\mathbb{C})$, which has a tracial state $\tau:=\tau_{0}\otimes\tau^{\prime}$ where $\tau^{\prime}$ is the unique tracial state on $M_{2(n+1)}(\mathbb{C})$. Finally,
@@
We now look at what happens when $m\not\in\mathcal{S}$. For all $n\geq 0$, let
@@
in *$\mathbb{C}^{*}\langle\mathcal{X}_{H}\cup\mathcal{X}_{0}\rangle$. Note that $\widetilde{X}_{n}^{2}=1$ in $\mathbb{C}\mathbb{Z}_{2}^{\mathcal{X}_{H}\cup\mathcal{X}_{0}}$.
@@
There are constants $C$ and $k$ such that for any $m\in\mathbb{N}$ and $0\leq n<h(m)$, $\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}-\widetilde{P}_{n+1}$ is trivial in $\mathcal{A}_{\mathcal{S}}(m)=\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}_{H}\cup\mathcal{X}_{0}}/\langle\mathcal{R}_{m}\rangle$, and has an $\mathcal{R}_{m}$-decomposition in $\mathbb{C}\mathbb{Z}_{2}^{\mathcal{X}_{H}\cup\mathcal{X}_{0}}$ of size $\leq C\big{(}(n+1)m\big{)}^{k}$.
Although we won’t need this, the constants $C$ and $k$ can be computed from a description of the Turing machine $M_{\mathcal{S}}$ that recognizes $\mathcal{S}$.
@@
By relation (R6), $fg$ is trivial in $\mathcal{A}_{\mathcal{S}}(m)$ and has an $\mathcal{R}_{m}$-decomposition in $\mathcal{A}$ of size $46n+$ 4. To prove the proposition, we show that
@@
in $\mathcal{A}_{\mathcal{S}}(m)$, while keeping track of the size using Lemma 3.3. For the first step, since $U$ and $S$ commute and $U\widetilde{X}U^{*}=S\widetilde{X}S^{*}$, we see that $\widetilde{X}_{n}=S^{n}\widetilde{X}S^{-n}$ in $\mathcal{A}_{\mathcal{S}}(m)$. Similarly $U^{n}\widetilde{Z}U^{-n}=T^{n}\widetilde{Z}T^{-n}$ and $U$ commutes with $J$, so
@@
Now suppose that $0\leq n<h(m)$. From the presentation of $G_{\mathcal{S}}$, we see that $X_{mn}Z_{mn}X_{mn}Z_{mn}=J$, so $Q(1+JX_{mn}Z_{mn}X_{mn}Z_{mn})=Q(1+J^{2})=2Q$. The word $JX_{mn}Z_{mn}X_{mn}Z_{mn}$ has length $516m+16n+$ 5, so by Proposition 4.8, part (c) there are $r_{1},\ldots,r_{k}\in\mathcal{R}_{H}\cup\mathcal{R}_{H}^{-1}$ and $z_{1},\ldots,z_{k}\in\mathcal{F}_{\mathcal{X}_{H}}$ with $JX_{mn}Z_{mn}X_{mn}Z_{mn}=z_{1}r_{1}z_{1}^{-1}\cdots z_{k}r_{k}z_{k}^{-1}$ in $\mathcal{F}_{\mathcal{X}_{H}}$, $k\leq d_{\mathcal{S}}(16m+16n+5)$, and $|z_{i}|\leq d_{\mathcal{S}}(16m+16n+5)$ for all $1\leq i\leq k$. Replacing $r_{i}$ with $1$ from right to left, we see that $Q(1+JX_{mn}Z_{mn}X_{mn}Z_{mn})-2Q$ has an $\mathcal{R}_{m}$-decomposition of size $O(d_{\mathcal{S}}(16m+16n+5)^{2})=O(d_{\mathcal{S}}(m)^{2}+d_{\mathcal{S}}(n)^{2})$. From Lemma 3.3, part (c), we get that $fg-2fQ$ has an $\mathcal{R}_{m}$-decomposition of size $O(n^{3}+n^{2}+nm+d_{\mathcal{S}}(m)^{2}+d_{\mathcal{S}}(m)^{2})$.
@@
has an $\mathcal{R}_{m}$-decomposition of size $O(n^{3})$. Ultimately we conclude that $\widetilde{P}_{n}+\widetilde{X}_{n}\cdot\widetilde{P}_{n}\cdot\widetilde{X}_{n}-\widetilde{P}_{n+1}$ has an $\mathcal{R}_{m}$-decomposition of size $O(n^{3}+n^{2}+nm+d_{\mathcal{S}}(m)^{2}+d_{\mathcal{S}}(m)^{2})$. ∎
@@
Let $\mathcal{X}=\mathcal{X}_{H}\cup\mathcal{X}_{0}$, and let $\widetilde{P}_{n}$, $\widetilde{X}_{n}$, and $\mathcal{R}_{m}$ be as above. The relations in (R1) are independent of $m$, and the remaining relations are spelled out explicitly in (R2)-(R6), so we see that the mapping $m\mapsto\mathcal{R}_{m}$ is computable. Parts (a) and (d) of the theorem hold by construction, and part (b) follows from Proposition 4.11. Finally, since $[P,Q]=0$ in $\mathcal{A}_{\mathcal{S}}(m)$, Lemma 4.10 implies that there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{S}}(m)$ with $\tau(\widetilde{P}_{0}^{2})=\tau(PQ)>0$. ∎
@@
Given $\alpha\in\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$, determine if $\alpha$ is trace-positive in *$\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$, or if $\pi(\alpha)$ is not trace-positive in $\mathbb{C}\mathbb{Z}_{2}^{*N}$, promised that one of the two is the case.
@@
Suppose that $N\geq n_{0}$, and that the homomorphism $\pi$ in Theorem 5.1 factors through a ***-algebra $\mathcal{A}$, meaning that there are homomorphisms $\pi_{1}:\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle\to\mathcal{A}$ and $\pi_{2}:\mathcal{A}\to\mathbb{Q}\mathbb{Z}_{2}^{*N}$ such that $\pi=\pi_{2}\circ\pi_{1}$. Suppose also that we have some way of representing the elements of $\mathcal{A}$ as strings such that the homomorphism $\pi_{1}$ is computable. Then Theorem 5.1 has the following immediate corollary:
@@
Suppose $\alpha\in\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$. If $\alpha$ is trace-positive in *$\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$, then $\pi_{1}(\alpha)$ is trace-positive in $\mathcal{A}_{\mathbb{C}}$. If $\pi(\alpha)$ is not trace-positive in $\mathbb{C}\mathbb{Z}_{2}^{*N}$, then there is a tracial state $\tau$ on $\mathbb{C}\mathbb{Z}_{2}^{*N}$ such that $(\tau\circ\pi_{2})(\pi_{1}(\alpha))=\tau(\pi(\alpha))<0$, so $\pi_{1}(\alpha)$ is not trace-positive in $\mathcal{A}_{\mathbb{C}}$. Since $\pi_{1}$ is computable, the promise problem in Theorem 5.1 reduces to the problem of determining if $\alpha\in\mathcal{A}$ is trace-positive. ∎
Theorem 1.3 is a special case of 5.2 with $\mathcal{A}=\mathbb{Q}^{*}\{x_{1},\ldots,x_{N}\}$, $\mathbb{Q}\mathcal{F}_{N}$, or $\mathbb{Q}\mathbb{Z}_{M}^{*N}$ (for every $M\geq M^{\prime}$, there is a ***-homomorphism $\mathbb{Q}\mathbb{Z}_{M}^{*N}\rightarrow\mathbb{Q}\mathbb{Z}_{M^{\prime}}^{*N}$ sending the $i$th generator of $\mathbb{Q}\mathbb{Z}_{M}^{*N}$ to the $i$th generator of $\mathbb{Q}\mathbb{Z}_{M^{\prime}}^{*N}$). The corollary also applies to other algebras such as the algebra of contractions and the algebra of self-adjoint contractions, as well as free products of any of the algebras mentioned.
For the proof of Theorem 5.1, suppose $\mathcal{S}\subseteq\mathbb{N}$ is an RE set, and let $\mathcal{X}$, $\widetilde{P}_{n}$ and $\widetilde{X}_{n}$, $n\geq 0$, and $\mathcal{A}_{\mathcal{S}}(m)=\mathbb{C}^{*}\langle\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}:\mathcal{R}_{m}\rangle$, $m\geq 1$, be as in Theorem 4.1. Before proceeding to the proof, we need the following propositions.
@@
There are positive integers $\Lambda,k$ such that if $m\notin\mathcal{S}$ and $\tau$ is a tracial $(\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ for some $\epsilon\geq 0$, then $\lVert\widetilde{P}_{0}\rVert_{\tau}\leq\Lambda m^{k}\sqrt{\epsilon}$.
@@
Let $C$ and $k$ be the constants from Theorem 4.1, part (b). Suppose $m\notin\mathcal{S}$, $\epsilon\geq 0$, and $\tau$ is a tracial $(\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By Theorem 4.1, part (a), $\widetilde{X}_{n}$ is a unitary involution in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. Since $\tau$ is tracial, $\lVert\widetilde{P}_{n}\rVert_{\tau}=\lVert\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert_{\tau}$ for all $n$. Also $\widetilde{P}_{n}$ and $\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}$ are hermitian squares in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, so $\tau(\widetilde{P}_{n}\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n})=\tau(\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\widetilde{P}_{n})\geq 0$. Hence
@@
for $m\in\mathbb{N}$. There are positive integers $\widetilde{\Lambda},k^{\prime}$ such that if $m\notin\mathcal{S}$ and $\tau$ is a tracial $(\epsilon,\mathcal{W}_{m})$-state on *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$, then $\lVert\widetilde{P}_{0}\rVert_{\tau}\leq\widetilde{\Lambda}m^{k^{\prime}}\sqrt{\epsilon}$.
@@
Let $\Lambda,k$ be integers from Proposition 5.3, and let $D$ be the integer from Theorem 4.1, part (d). Suppose $m\notin\mathcal{S}$, $\epsilon\geq 0$, and $\tau$ is a tracial $(\epsilon,\mathcal{W}_{m})$-state on *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$. By Lemma 3.11, there is a tracial state $\widehat{\tau}$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that
@@
Let $n_{0}:=\lvert\mathcal{X}\rvert$, and suppose we are given an integer $N\geq n_{0}$. Pick an ordering on $\mathcal{X}$ so we can identify *$\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ with the subalgebra *$\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n_{0}}\rangle$ of *$\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$. We think of *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$ and $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ as subalgebras of *$\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$ and $\mathbb{C}\mathbb{Z}_{2}^{*N}$ similarly. Let $\widetilde{\Lambda}$, $k^{\prime}$, and $\mathcal{W}_{m}$ be as in Proposition 5.4. For any $m\in\mathbb{N}$, define a ***-polynomial

### 18 Mar 2026 v2

@@ 4.2. Construction of $\mathcal{A}_{\mathcal{L}}(m)$
and similarly $\widetilde{\pi}(\widetilde{Z})\widetilde{\pi}(Q)=\widetilde{\pi}(Q)\widetilde{\pi}(\widetilde{Z})=\widetilde{\pi}(Z_{m,0})\widetilde{\pi}(Q)$. For (R4), recall from the presentation of $G_{\mathcal{L}}$ that $\pi(X_{m,0})=\pi(X_{m,n+1})$ and $\pi(S)\pi(X_{m,i})\pi(S^{*})=\pi(X_{m,i+1})$ for all $i\in\mathbb{Z}$. Since $\zeta E_{i,i}\zeta^{-1}=E_{i+1,i+1}$ for all $i\in\mathbb{Z}_{n+1}$,
@@
and $P_{n+1}:=P_{0}$. Recall from the presentation of $G_{\mathcal{L}}$ that $J$ is a unitary of order 2 and is central. Hence in $\mathbb{C}H_{\mathcal{L}}$, $\tfrac{1-J}{2}$ and $\tfrac{1+J}{2}$ are projections, and $(\tfrac{1+J}{2})P_{i}=0$ for $0\leq i\leq n$. Also for all $0\leq i\leq n-1$, $X_{mi}Z_{mi}X_{mi}Z_{mi}=J$, $X_{mi}$ commutes with $P_{i+1}$, and $P_{i}=(\tfrac{1-Z_{mi}}{2})P_{i+1}$. Hence $1+JX_{mi}Z_{mi}X_{mi}Z_{mi}=2$, and
@@
It follows that $\pi\big((P_{i}+X_{mi}P_{i}X_{mi}-P_{i+1})(1+JX_{mi}Z_{mi}X_{mi}Z_{mi})\big)=0$ for all $0\leq i\leq n$, and hence $A_{1}=A_{2}=0$. We conclude that $\widetilde{\pi}(r)=0$ for all $r$ in (R0)-(R6). Hence $\widetilde{\pi}$ induces a representation of $\mathcal{A}_{\mathcal{L}}(m)$ on $\widetilde{\mathcal{H}}\oplus\widetilde{\mathcal{H}}$.
The image of $\widetilde{\pi}$ is contained in $M_{2}\big(\pi(\mathbb{C}H_{\mathcal{L}})\otimes\mathcal{B}(\mathcal{H}_{1})\big)$ This algebra is isomorphic to $\pi(\mathbb{C}H_{\mathcal{L}})\otimes M_{2(n+1)}(\mathbb{C})$, which has a tracial state $\tau:=\tau_{0}\otimes\tau^{\prime}$ where $\tau^{\prime}$ is the unique tracial state on $M_{2(n+1)}(\mathbb{C})$. Finally,
@@
We now look at what happens when $m\not\in\mathcal{L}$. For all $n\geq 0$, let
@@
in $\mathbb{C}^{*}\langle\mathcal{X}_{H}\cup\mathcal{X}_{0}\rangle$. Note that $\widetilde{X}_{n}^{2}=1$ in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}_{H}\cup\mathcal{X}_{0}}$.
@@
There are constants $C$ and $k$ such that for any $m\in\mathbb{N}$ and $0\leq n<h(m)$, $\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}-\widetilde{P}_{n+1}$ is trivial in $\mathcal{A}_{\mathcal{L}}(m)=\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}_{H}\cup\mathcal{X}_{0}}/\langle\mathcal{R}_{m}\rangle$, and has an $\mathcal{R}_{m}$-decomposition in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}_{H}\cup\mathcal{X}_{0}}$ of size $\leq C\big((n+1)m\big)^{k}$.
Although we won’t need this, the constants $C$ and $k$ can be computed from a description of the Turing machine $M_{\mathcal{L}}$ that recognizes $\mathcal{L}$.
@@
By relation (R6), $fg$ is trivial in $\mathcal{A}_{\mathcal{L}}(m)$ and has an $\mathcal{R}_{m}$-decomposition in $\mathcal{A}$ of size $6n+4$. To prove the proposition, we show that
@@
in $\mathcal{A}_{\mathcal{L}}(m)$, while keeping track of the size using Lemma 3.3. For the first step, since $U$ and $S$ commute and $U\widetilde{X}U^{*}=S\widetilde{X}S^{*}$, we see that $\widetilde{X}_{n}=S^{n}\widetilde{X}S^{-n}$ in $\mathcal{A}_{\mathcal{L}}(m)$. Similarly $U^{n}\widetilde{Z}U^{-n}=T^{n}\widetilde{Z}T^{-n}$ and $U$ commutes with $J$, so
@@
Now suppose that $0\leq n<h(m)$. From the presentation of $G_{\mathcal{L}}$, we see that $X_{mn}Z_{mn}X_{mn}Z_{mn}=J$, so $Q(1+JX_{mn}Z_{mn}X_{mn}Z_{mn})=Q(1+J^{2})=2Q$. The word $JX_{mn}Z_{mn}X_{mn}Z_{mn}$ has length $16m+16n+5$, so by Proposition 4.8, part (c) there are $r_{1},\ldots,r_{k}\in\mathcal{R}_{H}\cup\mathcal{R}_{H}^{-1}$ and $z_{1},\ldots,z_{k}\in\mathcal{F}_{\mathcal{X}_{H}}$ with $JX_{mn}Z_{mn}X_{mn}Z_{mn}=z_{1}r_{1}z_{1}^{-1}\cdots z_{k}r_{k}z_{k}^{-1}$ in $\mathcal{F}_{\mathcal{X}_{H}}$, $k\leq d_{\mathcal{L}}(16m+16n+5)$, and $|z_{i}|\leq d_{\mathcal{L}}(16m+16n+5)$ for all $1\leq i\leq k$. Replacing $r_{i}$ with $1$ from right to left, we see that $Q(1+JX_{mn}Z_{mn}X_{mn}Z_{mn})-2Q$ has an $\mathcal{R}_{m}$-decomposition of size $O(d_{\mathcal{L}}(16m+16n+5)^{2})=O(d_{\mathcal{L}}(m)^{2}+d_{\mathcal{L}}(n)^{2})$. From Lemma 3.3, part (c), we get that fg2fQfg-2fQ has an $\mathcal{R}_{m}$-decomposition of size $O(n^{3}+n^{2}+nm+d_{\mathcal{L}}(m)^{2}+d_{\mathcal{L}}(n)^{2})$.
@@
has an $\mathcal{R}_{m}$-decomposition of size $O(n^{3})$. Ultimately we conclude that $\widetilde{P}_{n}+\widetilde{X}_{n}\cdot\widetilde{P}_{n}\cdot\widetilde{X}_{n}-\widetilde{P}_{n+1}$ has an $\mathcal{R}_{m}$-decomposition of size $O(n^{3}+n^{2}+nm+d_{\mathcal{L}}(m)^{2}+d_{\mathcal{L}}(n)^{2})$. ∎
@@
Let $\mathcal{X}=\mathcal{X}_{H}\cup\mathcal{X}_{0}$, and let $\widetilde{P}_{n}$, $\widetilde{X}_{n}$, and $\mathcal{R}_{m}$ be as above. The relations in (R1) are independent of $m$, and the remaining relations are spelled out explicitly in (R2)-(R6), so we see that the mapping $m\mapsto\mathcal{R}_{m}$ is computable. Parts (a) and (d) of the theorem hold by construction, and part (b) follows from Proposition 4.11. Finally, since $[P,Q]=0$ in $\mathcal{A}_{\mathcal{L}}(m)$, Lemma 4.10 implies that there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{L}}(m)$ with $\tau(\widetilde{P}_{0}^{2})=\tau(PQ)>0$. ∎
@@
Given $\alpha\in\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$, determine if $\alpha$ is trace-positive in $\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$, or if $\pi(\alpha)$ is not trace-positive in $\mathbb{C}\mathbb{Z}_{2}^{*N}$, promised that one of the two is the case.
@@
Suppose that $N\geq n_{0}$, and that the homomorphism $\pi$ in Theorem 5.1 factors through a *-algebra $\mathcal{A}$, meaning that there are homomorphisms $\pi_{1}:\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle\to\mathcal{A}$ and $\pi_{2}:\mathcal{A}\to\mathbb{Q}\mathbb{Z}_{2}^{*N}$ such that $\pi=\pi_{2}\circ\pi_{1}$. Suppose also that we have some way of representing the elements of $\mathcal{A}$ as strings such that the homomorphism $\pi_{1}$ is computable. Then Theorem 5.1 has the following immediate corollary:
@@
Suppose $\alpha\in\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$. If $\alpha$ is trace-positive in $\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$, then $\pi_{1}(\alpha)$ is trace-positive in $\mathcal{A}_{\mathbb{C}}$. If $\pi(\alpha)$ is not trace-positive in $\mathbb{C}\mathbb{Z}_{2}^{*N}$, then there is a tracial state $\tau$ on $\mathbb{C}\mathbb{Z}_{2}^{*N}$ such that $(\tau\circ\pi_{2})(\pi_{1}(\alpha))=\tau(\pi(\alpha))<0$, so $\pi_{1}(\alpha)$ is not trace-positive in $\mathcal{A}_{\mathbb{C}}$. Since $\pi_{1}$ is computable, the promise problem in Theorem 5.1 reduces to the problem of determining if $\alpha\in\mathcal{A}$ is trace-positive. ∎
Theorem 1.3 is a special case of 5.2 with $\mathcal{A}=\mathbb{Q}^{*}\{x_{1},\ldots,x_{N}\}$, $\mathbb{Q}\mathcal{F}_{N}$, or $\mathbb{Q}\mathbb{Z}_{M}^{*N}$ (for every $M\geq M^{\prime}$, there is a *-homomorphism $\mathbb{Q}\mathbb{Z}_{M}^{*N}\rightarrow\mathbb{Q}\mathbb{Z}_{M^{\prime}}^{*N}$ sending the $i$th generator of $\mathbb{Q}\mathbb{Z}_{M}^{*N}$ to the $i$th generator of $\mathbb{Q}\mathbb{Z}_{M^{\prime}}^{*N}$). The corollary also applies to other algebras such as the algebra of contractions and the algebra of self-adjoint contractions, as well as free products of any of the algebras mentioned.
For the proof of Theorem 5.1, suppose $\mathcal{L}\subseteq\mathbb{N}$ is an RE set, and let $\mathcal{X}$, $\widetilde{P}_{n}$ and $\widetilde{X}_{n}$, $n\geq 0$, and $\mathcal{A}_{\mathcal{L}}(m)=\mathbb{C}^{*}\langle\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}:\mathcal{R}_{m}\rangle$, $m\geq 1$, be as in Theorem 4.1. Before proceeding to the proof, we need the following propositions.
@@
There are positive integers $\Lambda,k$ such that if $m\notin\mathcal{L}$ and $\tau$ is a tracial $(\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ for some $\epsilon\geq 0$, then $\lVert\widetilde{P}_{0}\rVert_{\tau}\leq\Lambda m^{k}\sqrt{\epsilon}$.
@@
Let $C$ and $k$ be the constants from Theorem 4.1, part (b). Suppose $m\notin\mathcal{L}$, $\epsilon\geq 0$, and $\tau$ is a tracial $(\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By Theorem 4.1, part (a), $\widetilde{X}_{n}$ is a unitary involution in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. Since $\tau$ is tracial, $\lVert\widetilde{P}_{n}\rVert_{\tau}=\lVert\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert_{\tau}$ for all $n$. Also $\widetilde{P}_{n}$ and $\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}$ are hermitian squares in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, so $\tau(\widetilde{P}_{n}\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n})=\tau(\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\widetilde{P}_{n})\geq 0$. Hence
@@
for $m\in\mathbb{N}$. There are positive integers $\widetilde{\Lambda},k^{\prime}$ such that if $m\notin\mathcal{L}$ and $\tau$ is a tracial $(\epsilon,\mathcal{W}_{m})$-state on $\mathbb{C}^{*}\langle\mathcal{X}\rangle$, then $\lVert\widetilde{P}_{0}\rVert_{\tau}\leq\widetilde{\Lambda}m^{k^{\prime}}\sqrt{\epsilon}$.
@@
Let $\Lambda,k$ be integers from Proposition 5.3, and let $D$ be the integer from Theorem 4.1, part (d). Suppose $m\notin\mathcal{L}$, $\epsilon\geq 0$, and $\tau$ is a tracial $(\epsilon,\mathcal{W}_{m})$-state on $\mathbb{C}^{*}\langle\mathcal{X}\rangle$. By Lemma 3.11, there is a tracial state $\widehat{\tau}$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that
@@
Let $n_{0}:=\lvert\mathcal{X}\rvert$, and suppose we are given an integer $N\geq n_{0}$. Pick an ordering on $\mathcal{X}$ so we can identify $\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ with the subalgebra $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n_{0}}\rangle$ of $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$. We think of $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ and $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ as subalgebras of $\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$ and $\mathbb{C}\mathbb{Z}_{2}^{*N}$ similarly. Let $\widetilde{\Lambda}$, $k^{\prime}$, and $\mathcal{W}_{m}$ be as in Proposition 5.4. For any $m\in\mathbb{N}$, define a *-polynomial